find greatest and least value of fxn


f(x)= x+ 1/x


where a) x>0    (b) x belongs to R-{0}

Grade:12

1 Answers

Askiitians Expert Soumyajit IIT-Kharagpur
14 years ago

Dear STUDENT NAME,

Ans:- The given function is f(x)=x+1/x the 1st order derivative is=1-1/x²

The 2nd order derivative is=2/x³ Putting dy/dx=0 we get, x=+1, -1 Again checking the conditions for max && min in the 2nd order derivative we get the values as

Max value= -2

Min value=+2

Look carefully that the max value is less than that of the min value.

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best STUDENT NAME !!!

 


Regards,

Askiitians Experts
Soumyajit Das IIT Kharagpur